Using Butterworth responses, design active filters to meet the following specifications:2. A highpass filter with a 3dB frequency of 2kHz and 20dB attenuation at 1kHz.
Answer #1

2. Given that dp = 3dB at fp=2kHz. 2 2.2008 at fs=1k the e p = pau bond attenuation ds = stop band at tenutation to - pass ba2. Given that dp = 3dB at fp=2kHz. 2 2.2008 at fs=1k the e p = pau bond attenuation ds = stop band at tenutation to - pass baN 0.9988 -3.315 [ N = 9 : Tu ade a Bulte voorth, filter is q! List at Butter worth Polynomials INI Polynial of 14 (6) sti 5и лh CORP-1) 12566-37 (100 1x3 1) 18 de = 12573.83 radise Now substitute - - in equation 2 re HAS) = (C) +076596.) 17 6 5 [5+